# Analytic Scalar Field Theory

### v4

@@ Abstract
As an illustration of the method, consider the $O\left(\kappa\right)$ contribution to the vacuum expectation value (VEV) of the potential energy density. To this order in $\kappa$, only the free-field VEV of a Gaussian is required to complete the calculation. This is easily shown to be
@@
either by using operator methods or else by standard Feynman diagram combinatorics. The calculation therefore reduces to the free-field evaluation of $\left\langle\phi^{2}\right\rangle$, which can be done in ND using a momentum cut-off, or $\zeta$ function regularization, or dimensional regularization. All that remains is evaluation of the manifestly convergent contour integral in (2). Thus, by rescaling the variable of integration,
@@
So, at least to this order in $\kappa$, quantum effects have modified the form expected from classical field theory by a simple replacement of $\kappa$ with
@@
For example, the order $\kappa^{2}$ computation of $\left\langle V\left(\phi\right)\right\rangle$ follows from $\left\langle\left(\phi^{2}\left(0\right)\right)^{\nu}\left(\phi^{2}\left(x\right)\right)^{\nu}\right\rangle$ upon replacing the powers of $\phi^{2}$ with superpositions of Gaussians, as defined in (2). Evaluation of the VEV is then facilitated by the identity
@@
The generating functional for the latter VEVs is then obtained by taking $\sum_{n=0}^{\infty}\frac{\left(-z\right)^{n}}{n!}\phi^{2n}\left(x\right)$ leading to (9).
Once again for this example the integration over Hankel contours yields to routine evaluation methods. The final result is a Gauss hypergeometric function.

### v5

@@ Abstract
As an illustration of the method, consider the $O\left(\kappa\right)$ contribution to the vacuum expectation value (VEV) of the potential energy density. To this order in $\kappa$, only the free-field VEV of a Gaussian is required to complete the calculation. As shown in the Appendix this is
@@
The calculation therefore reduces to the free-field evaluation of $\left\langle\phi^{2}\right\rangle$, which can be done in ND using a momentum cut-off, or $\zeta$ function regularization, or dimensional regularization. All that remains is evaluation of the manifestly convergent contour integral in (2). Thus, by rescaling the variable of integration,
@@
So, at least to this order in $\kappa$, quantum effects have only modified the form expected from classical field theory by a simple replacement of $\kappa$ with
@@
For example, the order $\kappa^{2}$ computation of $\left\langle V\left(\phi\right)\right\rangle$ follows from $\left\langle\left(\phi^{2}\left(0\right)\right)^{\nu}\left(\phi^{2}\left(x\right)\right)^{\nu}\right\rangle$ upon replacing the powers of $\phi^{2}$ with superpositions of Gaussians as defined in (2). Evaluation of the VEV is then facilitated by the identity
@@
The generating functional for the latter VEVs is then obtained by taking $\sum_{n=0}^{\infty}\frac{\left(-z\right)^{n}}{n!}\phi^{2n}\left(x\right)$ to obtain (9).
For this example the integration over Hankel contours once again yields to routine evaluation methods. The final result is a Gauss hypergeometric function.
@@
Appendix: Some combinatorics
The free field VEV of a Gaussian follows from
$\left\langle\exp\left(-z\phi^{2}\right)\right\rangle=\sum_{n=0}^{\infty}\frac{\left(-z\right)^{n}}{n!}\left\langle\phi^{2n}\right\rangle$ (A1)
where $\left\langle\phi^{2n}\right\rangle$ is a “flower” graph with $n$ “petals” as sketched below, each petal representing the VEV $\left\langle\phi^{2}\right\rangle$.
A Feynman flower diagram.
The combinatoric factor for this graph is $\left(2n-1\right)!!=\left(2n\right)!/\left(2^{n}n!\right)$ where $\left(2n\right)!$ corresponds to the number of ways to select each individual $\phi$ from $\phi^{2n}$, each petal has a symmetry factor of $1/2$, and the additional $1/n!$ is a symmetry factor that corresponds to the number of permutations of the petals. That is to say,
$\left\langle\phi^{2n}\right\rangle=\left(2n-1\right)!!\left\langle\phi^{2}\right\rangle^{n}$ (A2)
This immediately gives the result (3) in the text, namely,
$\left\langle\exp\left(-z\phi^{2}\right)\right\rangle=\sum_{n=0}^{\infty}\frac{\left(2n-1\right)!!}{n!}\left(-z\left\langle\phi^{2}\right\rangle\right)^{n}=\frac{1}{\sqrt{1+2z\left\langle\phi^{2}\right\rangle}}$ (A3)
This result also follows from standard operator methods, e.g., as often used to prove the BCH formula.
As explained in the text following (9), similar combinatorics, as well as other methods, give the result for the VEV of two or more Gaussians.